OpenStudy (anonymous):

What value of b makes the statement true? 17 = b – (–6.2) A. –23.2 B. –10.8 C. 10.8 D. 23.2

OpenStudy (anonymous):

Just plug it in and see.

OpenStudy (anonymous):

17 = b - (-6.2) --- a negative times a negative equals a positive 17 = b + 6.2 --- subtract 6.2 from both sides 17 - 6.2 = b 10.8 = b 10.8 would make the statement true

OpenStudy (anonymous):

I got 23.2 because I used a calculater @kelliegirl33